OpenStudy (anonymous):

How to solve clock problem such as: What is the angle between the hands of a clock at 2:25?

OpenStudy (anonymous):

So is it 162.5 degree?

OpenStudy (anonymous):

Also, can you tell me what the 0 with a dash inside it mean?

hero (hero):

Just make sure you double-check your work.

OpenStudy (anonymous):

\[\theta \] ^What does that mean? I saw it in the website, and I'm not sure what that's suppose to mean.

hero (hero):

\(\theta\) means "theta" which in general means "angle" and in this case, "the angle between hour hand and minute hand"

OpenStudy (anonymous):

Ohh. So, \[\theta hr\] means the hour's angle?

hero (hero):

\(\theta = \left|\dfrac{60(2) - 11(25)}{2}\right| = 77.5\)

OpenStudy (anonymous):

Oh. I'm wrong. =/

hero (hero):

All you need is this formula to figure it out: \(\theta = \left|\dfrac{60H - 11M}{2}\right|\) H = hour hand M = minute hand

hero (hero):

I should have just given you this link instead: http://en.wikipedia.org/wiki/Clock_angle_problem#Equation_for_the_angle_between_the_hands

OpenStudy (anonymous):

That's okay. Still the same link. >_>